OpenStudy (anonymous):

the midpoint of uv is (5 -11). The coordinates of one point are U ( 3,5 ), Find the coordinates of endpoint of v.

OpenStudy (anonymous):

let coordinates of v be (x,y) (3+x)/2=5 and (5+y)/2=-11 3+x=10 5+y=-22 x=7 y=-22-5 y=-27

OpenStudy (anonymous):

yes lol I got it right lol I just wanted to make sure before I submitted thnks